OpenStudy (oleg3321):

Choose the slope-intercept equation of the line that passes through the point shown and is perpendicular to the line shown.

OpenStudy (oleg3321):

y =1/3 x y = -1/3 x + 4 y = 3x + 20 y = -3x - 16

OpenStudy (oleg3321):

@triciaal

OpenStudy (oleg3321):

@aum

OpenStudy (oleg3321):

@ShadowLegendX

OpenStudy (oleg3321):

@Ashleyisakitty

OpenStudy (oleg3321):

@Fifciol

OpenStudy (oleg3321):

@experimentX

OpenStudy (oleg3321):

please help

OpenStudy (experimentx):

looks like you didn't copy the question well ... there are few things missing.

OpenStudy (oleg3321):

theres a graph it intersects through the points -2,0 and 0,6

OpenStudy (experimentx):

okay find slope ... first

OpenStudy (oleg3321):

can you just please give me the answer im in a hurry i really am

OpenStudy (experimentx):

ok ok

OpenStudy (oleg3321):

thanks

OpenStudy (oleg3321):

please faster

OpenStudy (oleg3321):

@ganeshie8

OpenStudy (oleg3321):

@ganeshie8

OpenStudy (oleg3321):

ya thats the graph

OpenStudy (oleg3321):

but whats the answer

OpenStudy (experimentx):

slope is 3

OpenStudy (oleg3321):

ya but look at my answers at the top which one is it a b c or d?

OpenStudy (experimentx):

none

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=plot+y%3D3x%2B6